OpenStudy (anonymous):

Solve for k. 9/2k-3 = 4/k+1 -3 21 -21

sam (.sam.):

(9)/(2k-3)=(4)/(k+1) 9*(k+1)=4*(2k-3) 9k+9=4*(2k-3) 9k+9=8k-12 9k=-21+8k k=-21

OpenStudy (kaederfds):

K = -21

OpenStudy (anonymous):

cross multiply